A uniform, solid sphere of radius 4.00 cm and mass 2.25 kg starts with a purely translational speed of 2.25 m/s at the top of an inclined plane. The surface of the incline is 1.75 m long, and is tilted at an angle of 33.0∘ with respect to the horizontal. Assuming the sphere rolls without slipping down the incline, calculate the sphere's final translational speed ?2 at the bottom of the ramp.
Answer #1

r = radius of the solid sphere = 4 cm = 0.04 m

m = mass of solid sphere = 2.25 kg

Moment of inertia of solid sphere is given as

I = (0.4) m r2

L = length of incline surface = 1.75 m

\theta = angle of incline = 33

h = height of incline = L Sin\theta = (1.75) Sin33 = 0.953 m

v1 = translational speed at the top = 2.25 m/s

v2 = final translational speed at the bottom of ramp = ?

w = final angular speed at the bottom = v2 /r

Using conservation of energy

initial kinetic energy at top + Potential energy at top = kinetic energy at bottom + rotational kinetic energy at bottom

(0.5) m v12 + mgh = (0.5) m v22 + (0.5) I w2

(0.5) m v12 + mgh = (0.5) m v22 + (0.5) (0.4) (m r2) (v2/r)2

(0.5) m v12 + mgh = (0.5) m v22 + (0.2) m v22

(0.5) m v12 + mgh = (0.7) m v22

(0.5) v12 + gh = (0.7) v22

inserting the values

(0.5) (2.25)2 + (9.8 x 0.953) = (0.7) v22

v2 = 4.12 m/s